My son was given an assignment that I cannot figure out how to help with. He is to write three word problems that contain the concept of prime factorization and then explain the solutions.
Any chance you could give us an example of a word problem that does that? I STINK at math and if I could get an idea of what the teacher expects, I could probably help him write three. His assignment is due tomorrow, and without help, he will get a zero...

Thanks!

Answer by ankor@dixie-net.com(12704)

Here's a problem: Let say you are on a construction project and on the 1st day
you bought a certain number of 2 by 12's and paid by check, $1843. On the 2nd day you bought a different number of 2 by 12's for the same price and paid by check, $1957. As luck would have it you lost the itemized receipts. The accountant tells you I have to know the number of items, and their price for you to get reimbursed. You think hard and although you can't remember the price, you remember that it was a whole $ amt and a prime number:
:
Let the item cost = c;
:
Let x = no. of boards bought the 1st day; y = no. of boards on the 2nd day
:
cx = 1843 and cy = 1957, they have a common factor
:
We know that if 1843 and 1957 have a common factor, the difference will have the
same factor. Finding the difference will give a much small number to deal with
1957 - 1843 = 114; find the prime factors of 114: 2, 57,; factor 57 to 3, 19
Thus if 1843 and 1957 have a common factor it has to be 2, 3, 19, checks reveal
that it's obviously not 2, 3 doesn't work either, but what about 19?
1843/19 = 97 boards on the 1st day and 1957/19 = 103 boards on the 2nd day at $19 a board. The bean counter is happy and impressed by your math skill, and offer you a better position with firm, which started you on the road to success all because you know how to prime factor. (I got carried away)